OpenStudy (anonymous):

Solve 2x^2 - 4x + 8 = 0

OpenStudy (solomonzelman):

divide by 2, and complete the square.

OpenStudy (jdoe0001):

I assume you have covered the quadratic formula already?

OpenStudy (arbershabani97):

2(x^2-2x+4)=0 \[x^2-2x+4\] use quadratic formula

OpenStudy (solomonzelman):

\[2x^2 - 4x + 8 = 0\]\[2(x^2-2x+4)=0\]\[x^2-2x+4=0\]\[x^2-2x=-4\]\[x^2-2x+1=-4+1\]\[x^2-2x+1=-3\]\[(x-1)^2=-3\]\[x-1=± \sqrt{-3}\]\[x=1± \sqrt{-3}\]

OpenStudy (solomonzelman):

my favorite method:)

OpenStudy (jdoe0001):

\(\bf 2x^2 - 4x + 8 = 0\implies 2(x^2-2x+4)=0\implies {\color{blue}{ 1}}x^2{\color{red}{ -2}}x{\color{green}{ +4}}=0 \\ \quad \\ \textit{quadratic formula}\\ x= \cfrac{ - {\color{red}{ b}} \pm \sqrt { {\color{red}{ b}}^2 -4{\color{blue}{ a}}{\color{green}{ c}}}}{2{\color{blue}{ a}}}\)

OpenStudy (anonymous):

uhhhhh

OpenStudy (anonymous):

Does x = 6?

OpenStudy (anonymous):

(1,6)

OpenStudy (jdoe0001):

hmm what did you get from the quadratic formula or also check SolomonZelman by completing the square

OpenStudy (anonymous):

Simplifying 2x2 + -4x + -8 = 0 Reorder the terms: -8 + -4x + 2x2 = 0 Solving -8 + -4x + 2x2 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '2'. 2(-4 + -2x + x2) = 0 Ignore the factor 2. Subproblem 1 Set the factor '(-4 + -2x + x2)' equal to zero and attempt to solve: Simplifying -4 + -2x + x2 = 0 Solving -4 + -2x + x2 = 0 Begin completing the square. Move the constant term to the right: Add '4' to each side of the equation. -4 + -2x + 4 + x2 = 0 + 4 Reorder the terms: -4 + 4 + -2x + x2 = 0 + 4 Combine like terms: -4 + 4 = 0 0 + -2x + x2 = 0 + 4 -2x + x2 = 0 + 4 Combine like terms: 0 + 4 = 4 -2x + x2 = 4 The x term is -2x. Take half its coefficient (-1). Square it (1) and add it to both sides. Add '1' to each side of the equation. -2x + 1 + x2 = 4 + 1 Reorder the terms: 1 + -2x + x2 = 4 + 1 Combine like terms: 4 + 1 = 5 1 + -2x + x2 = 5 Factor a perfect square on the left side: (x + -1)(x + -1) = 5 Calculate the square root of the right side: 2.236067978 Break this problem into two subproblems by setting (x + -1) equal to 2.236067978 and -2.236067978. Subproblem 1 x + -1 = 2.236067978 Simplifying x + -1 = 2.236067978 Reorder the terms: -1 + x = 2.236067978 Solving -1 + x = 2.236067978 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + x = 2.236067978 + 1 Combine like terms: -1 + 1 = 0 0 + x = 2.236067978 + 1 x = 2.236067978 + 1 Combine like terms: 2.236067978 + 1 = 3.236067978 x = 3.236067978 Simplifying x = 3.236067978 Subproblem 2 x + -1 = -2.236067978 Simplifying x + -1 = -2.236067978 Reorder the terms: -1 + x = -2.236067978 Solving -1 + x = -2.236067978 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + x = -2.236067978 + 1 Combine like terms: -1 + 1 = 0 0 + x = -2.236067978 + 1 x = -2.236067978 + 1 Combine like terms: -2.236067978 + 1 = -1.236067978 x = -1.236067978 Simplifying x = -1.236067978 Solution The solution to the problem is based on the solutions from the subproblems. x = {3.236067978, -1.236067978} Solution x = {3.236067978, -1.236067978}